OpenStudy (lilith_mirabella):

last one @BlazeRyder 5. Suppose a skydiver (mass = 75 kg) is falling toward the Earth. When the skydiver is 100 m above the Earth he is moving at 60 m/s. At this point calculate the skydiver’s……. a. (2 pts.) Gravitational potential energy b. (2 pts.) Kinetic energy c. (2 pts.) Total mechanical energy (The sum of kinetic and potential energy)

OpenStudy (blazeryder):

Alright, so it would be the same 2 formulas: \[P.E.=\frac{ 1 }{ 2 } \times m \times v ^{2}\] and \[P.E._{g}=m \times g \times h\] So from the question: m=75kg H=100m, and v=60m/s So tell me what you get for A and B when we can work on C.

OpenStudy (blazeryder):

*then we can work on c

OpenStudy (lilith_mirabella):

i don't know

OpenStudy (blazeryder):

\[PE=0.5 \times 75 \times 60^{2}\] \[PE _{g}=75 \times 9.8 \times 100\] put those into a calculator and tell me what you get.

OpenStudy (lilith_mirabella):

for the first i think it's 2250

OpenStudy (lilith_mirabella):

the second one is 73500

OpenStudy (lilith_mirabella):

i think that's what the second one is

OpenStudy (blazeryder):

The first one is way off. And the second one is SOOOOOO close! I will give you the second one: it is 73,575. Tell me what you did to get your answer for a.

OpenStudy (lilith_mirabella):

i mulitplied the numbers and that's just what i got

OpenStudy (blazeryder):

Do you know what the 2 above the 60 is? If not that is okay, I can tell you =)

OpenStudy (lilith_mirabella):

2x

OpenStudy (lilith_mirabella):

times 2

OpenStudy (blazeryder):

Not quite, that sign means 60 times 60. Knowing this would would be the answer?

OpenStudy (lilith_mirabella):

3600

OpenStudy (blazeryder):

Let me break this one down for you =) \[KE=0.5 \times 75 \times 60^{2}\]\[KE=0.5*75*3600\]\[KE=75 \times 1800\]\[KE=135,000\] So B would be 135,000

OpenStudy (blazeryder):

Opps, sorry the KE should be PE.

OpenStudy (lilith_mirabella):

if i were to c's answer what would i do with the 2 numbers. Add Subtract Multiply or Divide?

OpenStudy (blazeryder):

Now Total Mechanical Energy =PE +PEg So: \[ME=135,000 +73,575\]

OpenStudy (lilith_mirabella):

208,575?

OpenStudy (blazeryder):

Perfect!

OpenStudy (lilith_mirabella):

thank you

OpenStudy (blazeryder):

No problem! And just for the future, this deals with Physics so if you could post in Physics, that would be wonderful! Here is the link so you know: http://openstudy.com/study#/groups/Physics

OpenStudy (lilith_mirabella):

okay. i'll remember that